# SAAS: Solving Ability Amplification Strategy for Enhanced Mathematical Reasoning in Large Language Models

### v3

@@ Abstract
Equal Contribution Corresponding
@@ 1 Introduction
2 Related Work and Background
Prompting Approach
Fine-tuning Approach
Continued Pretraining Approach
Some researches Lewkowycz et al. (2022); Azerbayev et al. (2023) continually pretrain a base model to specialize in the mathematical reasoning. Minerva Lewkowycz et al. (2022) is a large language model pretrained on general natural language data and further trained on the scientific and mathematical data. Llemma Azerbayev et al. (2023) was also obtained through continued pretraining Code Llama Roziere et al. (2023) on their own collected data named Proof-Pile-2333https://huggingface.co/datasets/EleutherAI/proof-pile-2.
@@ 3 SAAS: Solving Ability Amplification Strategy
Figure 2: Overall procedure of the synthetic data generation.
@@ 4.1.1 Dataset Details
Specifically, we synthesize these datasets into Chain-of-Thought (CoT) and Program-of-Thought (PoT) rationales via various models (GPT, WizardMath Luo et al. (2023), ToRA Gou et al. (2023)). To generate diverse synthetic data, we adjust some hyperparameters such as temperature and top_p. Then, we select only the correct responses and eliminate similar ones among these correct responses as in Wang et al. (2022). The detailed descriptions of seed datasets are described in Appendix A. Table 1 provides the summary of our synthetic datasets for fine-tuning.
@@ 4.1.3 Model Details
To evaluate the effectiveness of our SAAS in RQ1, we compared it with several state-of-the-art competitors. They can be divided into the following two groups:
  • General models: GPT-4 Achiam et al. (2023), ChatGPT (gpt-3.5-turbo) OpenAI (2023), Claude-2 Anthropic (2023), PaLM-2 Anil et al. (2023), LLaMA-2 Touvron et al. (2023), Platypus-2 Lee et al. (2023), CodeLLaMA Roziere et al. (2023), SOLAR-1 Kim et al. (2023).
  • Mathematics domain-specific models: WizardMath Luo et al. (2023), MetaMath Yu et al. (2023), MulggleMath Li et al. (2023a), Toolformer Schick et al. (2023), MathCoder Wang et al. (2023), MammoTH Yue et al. (2023), ToRA Gou et al. (2023).
@@ 4.2.4 RQ4: Case Study
As depicted in Figure 4, CoT learning approach exhibited inaccuracies in arithmetic computations as well as deficiencies in mathematical reasoning. Conversely, PoT approach demonstrated precise calculations yet exhibited a critical deficiency in mathematical reasoning. As we expected, our SAAS exhibited precise computational accuracy along with enhanced mathematical reasoning capabilities (See the more detailed comments than the comments of PoT learning). Through this case study, we demonstrated the following three observations: i) only CoT learning approach leads to arithmetic calculation errors; ii) only PoT learning approach may result in a deficit of mathematical reasoning; iii) sequential learning that transitions from CoT learning to PoT learning help improve computational accuracy as well as mathematical reasoning.
@@ Acknowledgements
Appendix A Detailed Descriptions of Seed Datasets

### v4

@@ Abstract
Equal Contribution Corresponding Author
@@ 3.1 Experimental Settings
Figure 2: Overall procedure of the synthetic data generation.
@@ 3.1.1 Dataset Details
Specifically, we synthesize these datasets into Chain-of-Thought (CoT) and Program-of-Thought (PoT) rationales via various models (GPT, WizardMath Luo et al. (2023), ToRA Gou et al. (2023)). To generate diverse synthetic data, we adjust some hyperparameters such as temperature and top_p. Then, we select only the correct responses and eliminate similar ones among these correct responses as in Wang et al. (2022). The detailed descriptions of seed datasets are described in Appendix B. Table 1 provides the summary of our synthetic datasets for fine-tuning.
@@ 3.1.3 Model Details
To evaluate the effectiveness of our SAAS in RQ1, we compared it with several state-of-the-art competitors. These competitors are divided into two groups: general models and mathematics domain-specific models. The general models include GPT-4 Achiam et al. (2023), ChatGPT (gpt-3.5-turbo)OpenAI (2023), Claude-2Anthropic (2023), PaLM-2 Anil et al. (2023), LLaMA-2 Touvron et al. (2023), Platypus-2 Lee et al. (2023), CodeLLaMA Roziere et al. (2023), and SOLAR-1 Kim et al. (2023). The mathematics domain-specific models consist of WizardMath Luo et al. (2023), MetaMath Yu et al. (2023), MulggleMath Li et al. (2023a), Toolformer Schick et al. (2023), MathCoder Wang et al. (2023), MammoTH Yue et al. (2023), and ToRA Gou et al. (2023).
@@ 3.2.4 RQ4: Case Study
As depicted in Figure 4, CoT learning approach exhibited inaccuracies in arithmetic computations as well as deficiencies in mathematical reasoning. Conversely, PoT approach demonstrated precise calculations yet exhibited a critical deficiency in mathematical reasoning. As we expected, our SAAS exhibited precise computational accuracy along with enhanced mathematical reasoning capabilities (See the more detailed comments than the comments of PoT learning). Through this case study, we demonstrated the following three observations: i) only CoT learning approach leads to arithmetic calculation errors; ii) only PoT learning approach may result in a deficit of mathematical reasoning; iii) sequential learning that transitions from CoT to PoT learning help improve computational accuracy as well as mathematical reasoning.
@@ Acknowledgements
Appendix A Related Work and Background
Prompting Approach
Fine-tuning Approach
Continued Pretraining Approach
Some researches Lewkowycz et al. (2022); Azerbayev et al. (2023) continually pretrain a base model to specialize in the mathematical reasoning. Minerva Lewkowycz et al. (2022) is a large language model pretrained on general natural language data and further trained on the scientific and mathematical data. Llemma Azerbayev et al. (2023) was also obtained through continued pretraining Code Llama Roziere et al. (2023) on their own collected data named Proof-Pile-2555https://huggingface.co/datasets/EleutherAI/proof-pile-2.
Appendix B Detailed Descriptions of Seed Datasets